OpenStudy (stormoshawty):

During a clothing store's Bargain Days, the regular price for T-Shirts is discounted by $5. There is a state sales tax of 5%, and the $5 discount is applied before the sales tax is calculated. A. Write and expression that shows the regular price r of a T-Shirt minus the $5 discount. B. Write a rule for the function p(r) that expresses the final price p of a t-shirt with the discount applied and sales tax added. C. How much would you pay during Bargain Days for a shirt regularly priced at $15.50?

OpenStudy (shamil98):

A. , just the discount applied to the regular price so r-5 B. p(r) = (r-5) + 0.05(r-5) C. Evaluate it with the given formula p(15.5) = (15.5 - 5) + 0.05(15.5-5)

OpenStudy (shamil98):

To explain B better.. 5% of something is just 0.05 of that something 0.05(x) for this equation you are asked for x - 5 so it becomes 0.05(x-5)

OpenStudy (stormoshawty):

God, thank you so much, Lol. What would i do without you? xD

OpenStudy (shamil98):

lol, you're welcome

OpenStudy (shamil98):

p(15.5) = (15.5 - 5) + 0.05(15.5-5) p(15.5) = 10.5 + 0.05(10.5) p(15.5) = 10.5 + 0.0525 = 10.5525 = 10 dollars and 55 cents